OpenStudy (anonymous):

i tried and tried, can't figure it out...what is the maximum height of a cliff diver if he jumps from 48 feet at 8ft/sec?

OpenStudy (anonymous):

i'm supposed to use this equation h=-16t^2 + vt + s h : the vertical component of the diver’s motion v : The initial upward velocity s : The height of the ledge above the water t : the time in seconds

OpenStudy (anonymous):

help please, its due tomorrow...

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

its not funny...

jimthompson5910 (jim_thompson5910):

in this case v = 8 and s = 48 so h=-16t^2 + vt + s turns into h=-16t^2 + 8t + 48

jimthompson5910 (jim_thompson5910):

h=-16t^2 + 8t + 48 graphs out a parabola that opens downward The vertex is the point where the max height is reached Use this formula to find the vertex t = -b/(2a) t = -8/(2*(-16)) ... plug in a = -16 and b = 8 t = -8/(-32) t = 1/4 t = 0.25 This means that at t = 0.25 seconds, the person is at max height

OpenStudy (anonymous):

ok, then i plug it in, right?

OpenStudy (anonymous):

its 49?

jimthompson5910 (jim_thompson5910):

Plug in t = 0.25 to get h=-16t^2 + 8t + 48 h=-16(0.25)^2 + 8(0.25) + 48 h=-16(0.0625) + 8(0.25) + 48 h=-1 + 2 + 48 h = 49 So the person is at the max height of 49 ft (at the vertex) when the time is t = 0.25 seconds

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

np